# SciEx: Benchmarking Large Language Models on Scientific Exams with Human Expert Grading and Automatic Grading

### v1

@@ 4.3.3 Influential Factors
Table 5: Graders performance on different examinees.
@@
Table 6: Graders performance on difficulty levels.
@@ 5 Conclusion
There are certain biases that can occur for SciEx. Firstly, the LLMs do not have time pressure. Therefore, they can output longer answers, which helps them get better grades, as there is a higher likelihood that something will be correct. Secondly, the grading process can not be fully anonymized. It is not easy to mix the LLM answers with student answers for the lecturer to grade, since student answers are usually handwritten. Therefore, the lecturers know when they are grading an LLM, thus can bias the score they give. Thirdly, the comparison between the LLMs and the students might be unfair, since the students studied the centralized course material specifically for the exams, while this is not the case for the LLMs. Lastly, due to the reliance on expert resources, the size of SciEx is quite small compared to other scientific benchmarks.
@@
For automatic grading, regardless of the high correlation to expert grading, they can still be imperfect. We are not suggesting to use LLMs to evaluate students, but to evaluate new models coming out when it is not possible to do human evaluation.
@@ Appendix B Exam Formatting
Originally, exams were in different formats, depending on their creator. We convert the exams into the JSON format, with file paths to images if any. An example is shown in Figure 4.
@@
Appendix C User Interface for Expert Grading
@@
Appendix D Performance on Text-only, English-only Questions
@@
Appendix E Grader Performance
E.1 Pearson Correlation on Exam Level
@@
E.2 RMSE on Question-level
@@
Table 9: LLM grading’s RMSE compared to expert grading on the question level. Note that there are only a single scores for zero-shot, since they do not have different shot settings.
E.3 Performance on Giving Full Points

### v3

@@ 4.3.3 Influential Factors
Table 6: LLM graders performance (i.e., Pearson correlation to expert grading) on different examinees.
@@
Table 7: LLM grader’s performance (i.e., Pearson correlation to expert grading) on different difficulty levels.
@@ 5 Conclusion
There are certain biases that can occur for SciEx. Firstly, the LLMs do not have time pressure. Therefore, they can output longer answers, which helps them get better grades, as there is a higher likelihood that something will be correct. Secondly, the grading process can not be fully anonymized. It is not easy to mix the LLM answers with student answers for the lecturer to grade, since student answers are usually handwritten. Additionally, the LLMs’ answers content itself might also be easily distinguishable from the students’, since the LLMs tend to, e.g., give longer answers or repeat the questions. Therefore, the lecturers know when they are grading an LLM, thus can bias the score they give. Thirdly, the comparison between the LLMs and the students might be unfair, since the students studied the centralized course material specifically for the exams, while this is not the case for the LLMs. Lastly, due to the reliance on expert resources, the size of SciEx is quite small compared to other scientific benchmarks.
@@
Automatic grading, regardless of the high correlation to expert grading, can still be imperfect. We are not suggesting to use LLMs to evaluate students, but to evaluate new models coming out when it is not possible to do human evaluation.
Regarding data consent, we had group meetings and email exchanges to come to an agreement from all lecturers that the data would be made public under the CC BY-NC-SA 4.0 license.
@@ Acknowledgments
This work was supported by the Helmholtz Programme-oriented Funding, with project number 46.24.01, project name AI for Language Technologies. It was also supported by funding from the pilot program Core-Informatics of the Helmholtz Association (HGF).
We thank the lecturers for their contribution during the creation of the dataset: Kunyu Peng, Alexander Jaus, David Schneider, Ruiping Liu, Zdravko Marinov, Yufan Chen, Miklós Borsi, Florian Kalinke, Federico Matteucci, Fabian Richter, Bela Böhnke, Jose Cribeiro-Ramallo, Daniel Ebi, Florian Kalinke, Adrian Feilhauer, Wendy Yi, Laura Merker, Miriam Goetze, Jean-Pierre von der Heydt, Max Göttlicher, Thomas Bläsius, Marcus Wilhelm, Michael Zündorf
@@ Appendix B Exam Formatting
Originally, exams were in different formats, depending on their creator. We convert the exams into JSON format, with file paths to images if any. An example is shown in Figure 4.
@@
Appendix C LLM Answer Generation Prompts
We provide the prompt in the same language as the exam question to the LLMs to generate answers. The English prompt is shown in Figure 5, and the German prompt is shown in Figure 6.
You are a university student. Please answer the following JSON-formatted exam question. The subquestions (if any) are indexed. The provided figures (if any) each contains its path at the bottom, which matches the path provided in the JSON. Please give the answers to the question and subquestions that were asked, and index them accordingly in your output. You do not have to provide your output in the JSON format. If you are asked to draw on the figure, then describe with words how you would draw it. Please provide all answers in English. Here is the question: <input text>
Figure 5: Answer generation prompt in English.
Sie sind Student. Bitte beantworten Sie die folgende JSON-formatierte Prüfungsfrage. Die Unterfragen (falls vorhanden) sind indiziert. Die bereitgestellten Abbildungen (falls vorhanden) enthalten jeweils unten ihren Pfad, der mit dem im JSON bereitgestellten Pfad übereinstimmt. Bitte geben Sie die Antworten auf die gestellten Fragen und Unterfragen an und indizieren Sie diese in Ihrer Ausgabe entsprechend. Sie müssen Ihre Ausgabe nicht im JSON-Format bereitstellen. Wenn Sie aufgefordert werden, auf der Figur zu zeichnen, beschreiben Sie mit Worten, wie Sie sie zeichnen würden. Bitte geben Sie alle Antworten auf Deutsch an. Hier ist die Frage: <input text>
Figure 6: Answer generation prompt in German.
Appendix D LLM Grader Prompts
We provide the prompt in the same language as the exam question to the LLMs to perform automatic grading. The English prompt is shown in Figure 7, and the German prompt is shown in Figure 8.
You are a university professor. Please grade the following exam question. The exam question, examinee’s answer, correct answer, and the maximum possible score are provided in the format: [question] <exam_question> [/question] [answer] <answer> [/answer] [correct_answer] <correct_answer> [/correct_answer] [max_score] <max_score> [/max_score] The question is provided in JSON format, but the answer can be freeform text. The provided figures in the question (if any) each contain its path at the bottom, which matches the path provided in the JSON. The answer is text-only. If the question asks to draw on the figure, then the answer should contain a text description of how the drawing should be. Please provide the grade between [0, <max_score>]. Please provide the reasoning for your grade. Please provide your output in the format: [reason] <reasoning> [/reason] [grade] <grade> [/grade] Below you are provided with examples on how to perform the grading: <example text> Here is your input: <input text>
Figure 7: Grading prompt in English.
Sie sind Universitätsprofessor. Bitte bewerten Sie die folgende Prüfungsfrage. Die Prüfungsfrage, die Antwort des Prüflings, die richtige Antwort und die maximal mögliche Punktzahl werden im Format bereitgestellt:: [question] <Prüfungsfrage> [/question] [answer] <Antwort> [/answer] [correct_answer] <korrekteAntwort> [/correct_answer] [max_score] <maxPunkt> [/max_score] Die Frage wird im JSON-Format bereitgestellt, die Antwort kann jedoch Freiformtext sein. Die bereitgestellten Abbildungen in der Frage (falls vorhanden) enthalten jeweils unten ihren Pfad, der mit dem im JSON bereitgestellten Pfad übereinstimmt. Die Antwort ist nur Text. Wenn es sich bei der Frage darum handelt, auf der Abbildung zu zeichnen, sollte die Antwort eine Textbeschreibung darüber enthalten, wie die Zeichnung aussehen soll. Bitte geben Sie die Note zwischen [0, <maxPunkt>] an. Bitte begründen Sie Ihre Note. Bitte geben Sie Ihre Ausgabe im Format an: [reason] <Grundsatz> [/reason] [grade] <Note> [/grade] Nachfolgend finden Sie ein Beispiel für die Durchführung der Benotung: <example text> Hier ist Ihre Eingabe: <input text>
Figure 8: Grading prompt in German.
Appendix E User Interface for Expert Grading
@@
Appendix F Performance on Text-only, English-only Questions
@@
Appendix G Grader Performance
G.1 Pearson Correlation on Exam Level
@@
G.2 RMSE on Question-level
@@
Table 10: LLM grading’s RMSE compared to expert grading on the question level. Note that there are only single scores for zero-shot, since they do not have different shot settings.
G.3 Performance on Giving Full Points